deternine the resultant of the following coplanar forces. 15N at 30 degress from the horizaontal, 55N at 80 degrees, 90N at 210 degrees, and 130N at 260 degrees


1
Expert's answer
2021-09-30T13:00:03-0400

"\vec F_1= 15N \ \angle(\vec F_1) =30"

"\vec F_2= 55N \ \angle(\vec F_2) =80"

"\vec F_3= 90N \ \angle(\vec F_3) =210"

"\vec F_4= 130N \ \angle(\vec F_4) =260"

"\vec F = \vec F_1+\vec F_2+\vec F_3+\vec F_4"

"\vec{F_5}=\vec{F_1}+\vec{F_3}\ \angle{(\vec{F_3},\vec{F_1})}=210-30=180"

"\vec F_5= 90-15 =75N;\angle(\vec F_5) =210"

"\vec{F_6}=\vec{F_2}+\vec{F_4}\ \angle{(\vec{F_3},\vec{F_1})}=260-80=180"

"\vec F_6= 130-55 =75N;\angle(\vec F_6) =260"

"\vec F = \vec F_5+\vec F_6"

"|\vec F_5|=|\vec F_6|"

"\angle (\vec F_6,\vec F_5)= 260-210 =50"

"\angle(\vec F) =\angle(\vec F_5)+\frac{\angle(\vec F_6)-\angle(\vec F_5)}{2}=235"

"|\vec F|^2= |\vec F_5|^2+ |\vec F_6|^2-2 |\vec F_5| |\vec F_6|\cos\angle (\vec F_6,\vec F_5)"

"|\vec F|^2= 75^2+75^2-2*75^2*\cos(50\degree)=4018.64"

"|\vec F|=\sqrt{4018.64}=63.39"

"\text{Answer: }\vec F=63.39N;\angle(\vec F) =235\degree"